solve by completing the square
m^2 – 1/2m = 7/2

Answer by mananth(16949)

m^2 – 1/2m = 7/2
m^2-1/2 m +1/16 = 7/2 +1/16
(m-1/4)^2=(56+1)/16
(m-1/4)^2=57/16
take the square root
(m-1/4)= +/- sqrt(57)/4
m=1/4+sqrt(57)/4 OR 1/4-sqrt(57)/4